OpenStudy (anonymous):

If both x and y are in quadrant IV, and sin x = -2/3 and cos y = 1/4. Find cos (x-y)

OpenStudy (anonymous):

cosxcosy+sinxsiny

OpenStudy (anonymous):

cos( x-y)= cos x cosy+ sinx siny

OpenStudy (anonymous):

sin x= -2/3 cox x =sqrt(1 +4/9)=sqrt(13)/3

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

cosx = sqrt(1-4/9)= sqrt(5)/3

OpenStudy (anonymous):

cosy =1/4 siny =- sqrt(1-1/16)=sqrt(15)/4

OpenStudy (anonymous):

put the values in formula

OpenStudy (anonymous):

ok thanks....another questions is to find sin (x+y) with the same variables... can you help me get started on that one also?

OpenStudy (anonymous):

oops, once would have been enough

OpenStudy (anonymous):

sin(x+y)= sinxcosy-cosxsiny

OpenStudy (anonymous):

values are same

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so for cos (x-y) would it be: (√5/3)(1/4)+(-2/3)(√15/4)=(√5/12)+(-2√15/12)= √5-2√15/12

OpenStudy (anonymous):

siny is negative in fourth quadrant sin y =-sqrt(15)/4

OpenStudy (anonymous):

ok so it would be √5+2√15/12

OpenStudy (anonymous):

for sin (x+y) would it be (-2/3)(1/4)-(√5/3)((√15/4)= -2/12-√75/12=-2-5√3/12 are my negatives and positives right?

OpenStudy (anonymous):

u again missed that siny is negative

OpenStudy (anonymous):

-2/12+√75/12

OpenStudy (anonymous):

for sin (x+y) would it be (-2/3)(1/4)-(√5/3)((√15/4)= -2/12-√75/12=-2-5√3/12 are my negatives and positives right?